# Operational research approaches and mathematical models for kidney exchange: A literature survey and empirical evaluation

### v1

@@ 2.8 Dataset Generators and Other Software Tools
Dataset generators and software tools play an important role in facilitating simulation studies that help to inform policy decision-making relating to KEPs. Often historical KEP medical data are either not available or not suitable for simulations, making dataset generators highly valuable to researchers and practitioners. Such tools can simulate real-world data under hypothetical changes to the recipient and donor pool, such as a larger number of NDDs and RDPs. These generators fall into two categories: static generators produce data for a single matching run, and dynamic generators simulate multiple consecutive matching runs as RDPs and NDDs arrive and depart from the pool. We also review software tools for KEPs. These can allow optimal solutions to be found for a given KEP instance under a range of different constraints and optimality objectives. Moreover, they can allow researchers and practitioners to better understand the long-term effects of different policies informing match runs.
@@
We first consider static dataset generators, the first of which was created by Saidman et al., (2006). Their generator takes into account factors such as ABO-compatibility, sensitisation and positive crossmatch probability to create a randomly generated pool of RDPs (note that they did not include NDDs). This generator was used in various subsequent simulations to model KEPs, for instance Roth et al., (2007); Abraham et al., (2007); Dickerson et al., (2013); Constantino et al., (2013); Ashlagi and Roth, (2021). There are 310 instances created by the Saidman generator that are publicly available on PrefLib (Mattei and Walsh, , 2013), initially generated by Dickerson et al., 2012b , some of which also include NDDs.
Delorme et al., (2022) explored differences between datasets produced by the Saidman generator (Saidman et al., , 2006) and historical data from the UK KEP. They found that the Saidman generator typically produces instances that are inconsistent with real UK KEP datasets with respect to a number of measures. They implemented a new dataset generator to produce instances reflecting key characteristics of UK KEP data; also, their generator can include NDDs.
Two additional static generators were created by Dickerson et al., 2012b and Ashlagi et al., (2013). The former altered the Saidman generator to create sparser instances that reflect the UNOS pool (UNOS being a leading US KEP), whilst the latter created only two types of recipients within their instances: those with low and high PRA [^2] PRA, which stands for Panel Reactive Antibody, measures the sensitisation of a recipient, and indicates the likelihood of the recipient being incompatible with a random blood-group compatible donor from the population. These generators were used by Arslan et al., (2024) in addition to the aforementioned Saidman instances from PrefLib.
@@
Dynamic generators allow users to analyse the long-term evolution of a KEP over multiple matching runs. Santos et al., (2017) created a modular and configurable dynamic dataset generator and KEP simulator that allows KEP pools to be created with diverse characteristics, allowing for incompatible RDPs, recipients with multiple donors, compatible RDPs and NDDs. Their tool includes configuration, pool management and optimisation modules and allows the simulation to be controlled by various parameters. The authors also compared their generation and simulation tool with various other dynamic generators in the literature. A more recent dynamic simulator, known as the ENCKEP Simulator, was developed as part of the ENCKEP COST Action (ENCKEP, , 2021). Its optimisation engine allows a range of constraints and optimality objectives to be specified and evaluated in a simulated KEP. It was extended and used in simulations by Matyasi and Biró, (2023) and Druzsin et al., (2024). The former advanced the ENCKEP simulator to test re-optimisation strategies, whilst the latter expanded it to allow for various optimisation criteria and evaluated these criteria in an international setting. Furthermore, Carvalho et al., (2024) used dynamic simulations based on the Canadian Kidney Paired Donation Programme to learn optimal weights to associate with each RDP and NDD used in their selection method. Others have used dynamic dataset generators to model the characteristics of programmes in the USA, such as Sönmez et al., (2020).
@@
Here we mention some additional resources, including web applications and visualisers for kidney exchange. Examples of such applications are linked to from https://www.dcs.gla.ac.uk/~davidm/software.html, including (i) a kidney exchange “toolkit” that can solve KE-Opt instances under different hierarchical optimality objectives, written by James Trimble; (ii) a KEP static dataset generator written by William Pettersson and James Trimble, (iii) a tool to visualise the compatibility graph and an optimal solution for a KE-Opt instance written by John Debois; and (iv) an interactive kidney exchange game for outreach and public engagement, written by William Pettersson.
@@ 2.9 Emerging Topics
Recent medical advancements have enabled the possibility of transplants between donors and recipients that are ABO- or HLA-incompatible (Andersson and Kratz, , 2020; MacMillan et al., , 2023). However, these transplants can be expensive, resource-intensive and recipients may have to take immunosuppressant medication post-transplant (see, e.g., the overviews from Montgomery et al., 2011 and Colaneri, 2014). Aziz et al., (2021) modelled KE-Opt in the presence of half-compatible arcs, which represent arcs $(v_{1},v_{2})$ in the underlying compatibility graph where the recipient of $v_{2}$ is not immediately compatible with the donor of $v_{1}$, but can become compatible with this donor through a medical procedure or treatment, such as the use of immunosuppressants. They studied the problem of finding an optimal set of exchanges given a budget that limits the number of half-compatible arcs that can be used. The special case of the problem in which all arcs are either compatible or half-compatible was studied by Heo et al., (2021) from a mechanism design point of view and by Delorme et al., (2025) from a modelling point of view.
@@
As mentioned in Section 1, KEPs may include compatible RDPs, where their participation could give a better match for the recipient compared to their willing donor, and could benefit the remaining KEP pool by giving additional options, especially for hard-to-match recipients (Gentry et al., , 2007). Li et al., (2019) presented a new algorithm and conducted simulations for KEPs with compatible RDPs, taking into account the fact that compatible RDPs will typically expect to be matched quickly (or else they would leave the pool and match directly). Sönmez et al., (2020) proposed incentivising compatible RDPs to join a KEP by providing the recipient with a priority in the DDWL, should they require a repeat transplant. Finally, we mention the work of Balbuzanov, (2020), who studied incentivising compatible pairs to participate, in a setting with ordinal recipient preferences, considering criteria such as individual rationality and Pareto optimality.
@@
Another initiative, namely Global Kidney Exchange (GKE), has recently been the focus of some attention in the literature since it was first proposed by Rees et al., (2017). GKE involves finding an RDP $v_{1}$ from a country $A$, where the recipient of $v_{1}$ is compatible with the donor of an RDP $v_{2}$ in a country $B$. Typically $A$ is a developing country without its own KEP, where, even if $v_{1}$ is a compatible RDP, they are financially incompatible, meaning that $v_{1}$ cannot afford the cost of the surgery. Moreover, $B$ is generally a developed country with its own KEP, such as the US, and $v_{2}$ may involve a hard-to-match recipient. Pair $v_{1}$ travels to country $B$ for the surgery, enabling the recipient of $v_{2}$ to obtain a transplant; and post-transplant care for pair $v_{1}$ in country $A$ is paid for by country $B$. Although GKE can provide additional transplantation opportunities, the ethics of this form of kidney exchange have been widely debated (see, e.g., Minerva et al., 2019; Ambagtsheer et al., 2020).
@@ 2.10 Other Related Surveys
There are a range of surveys that emphasise different aspects of KEPs. For example, Mak-Hau, (2017) gave a comprehensive survey of the state of the art in terms of ILP models for kidney exchange at the time of writing. Ashlagi and Roth, (2021) also focussed on summarising KEPs from an operational point of view, specifically in terms of what makes a KEP successful, and they also outlined a range of directions for future research in relation to KEPs. Broad historical overviews of the development of KEPs are given by Gentry et al., (2011), Glorie et al., 2014a and Kher and Jha, (2020), while Sönmez and Ünver, (2013); Ashlagi, (2023); Sönmez and Ünver, (2025) survey KEPs from a market design perspective. The survey of van Basshuysen, (2020) was given from a philosophical and ethical standpoint. Viana et al., (2022) discussed the complexity of KE-Opt, presenting models and algorithms, and also covered stochastic and robust models. Doval, (2025) provided a detailed survey of dynamic matching, making reference to dynamic KEPs. We also mention the short recent survey given by Sharifi, (2025).
Several papers presented descriptions of the national KEPs in various countries, including Australia (Cantwell et al., , 2015; Sypek et al., , 2017), Canada (Malik and Cole, , 2014), France (Combe et al., , 2019), India (Kute et al., , 2021), Spain (Bofill et al., , 2017), the UK (Johnson et al., , 2008) and the USA (Agarwal et al., , 2018; Flechner et al., , 2018). When focusing on KEPs across Europe, differences in their medical practices were surveyed by Biró et al., 2019a , while differences in their modelling and optimisation were surveyed by Biró et al., (2021).
@@ 3 A Detailed Exposition of ILP Formulations for KE-Opt
Whilst some algorithms for KE-Opt were surveyed in Section 2.1, we deferred our detailed coverage of ILP-based approaches to this section. We focus on the fundamental version of KE-Opt, as introduced in Section 1, excluding elements such as hierarchical objectives, deceased-donor-initiated chains, and other extensions discussed in Section 2. However, we emphasize that ILP models incorporating these advanced extensions are typically built upon the formulations covered here.
This section is structured as follows. In Section 3.1 we lay the foundations for our formal descriptions of a range of ILP models by providing a formal definition of KE-Opt, together with associated notation and terminology. Next, in Section 3.2 we introduce the ILP models for the version of KE-Opt in which only cycles are considered, after which we explain that these models can be adapted to instead model a version of KE-Opt with only chains. This is followed by a discussion on how to model the cycles-and-chains version of KE-Opt. Finally, in Sections 3.33.7 we present the ILP models in detail.
@@ 3.1 Formal Problem Statement
An instance of KE-Opt involves the set $\mathcal{R}$ of recipient-donor-pairs (RDPs), the set $\mathcal{N}$ of non-directed donors (NDDs) and the directed compatibility graph $\mathcal{G}=(\mathcal{V},\mathcal{A})$, as well as the limit $K$ on the maximum cycle length and the limit $L$ on the maximum chain length. Vertex set $\mathcal{V}=\mathcal{R}\cup\mathcal{N}\cup\{\tau\}$ contains a node for every RDP in $\mathcal{R}$ and for every NDD in $\mathcal{N}$, and also a terminal node $\tau$ that represents either the Deceased Donor Waiting List (DDWL) or the pool of bridge donors for subsequent matching rounds. Arc set $\mathcal{A}=\mathcal{A}_{\mathcal{R}}\cup\mathcal{A}_{\mathcal{N}}\cup\mathcal{A}_{\tau}$ contains three types of arcs. First, there is an arc $(u,v)\in\mathcal{A}_{\mathcal{R}}$ for each $u,v\in\mathcal{R}$ such that the donor of RDP $u$ is compatible with the recipient of RDP $v$. Second, there is an arc $(u,v)\in\mathcal{A}_{\mathcal{N}}$ for each $u\in\mathcal{N}$ and $v\in\mathcal{R}$ such that NDD $u$ is compatible with the recipient of RDP $v$. Third, there is an arc $(v,\tau)\in\mathcal{A}_{\tau}$ from every RDP or NDD $v\in\mathcal{R}\cup\mathcal{N}$ to the terminal node $\tau$.
@@
This definition of $\mathcal{G}$ allows for modelling any sequence of exchanges by either a cycle or a chain. A cycle of length $k$, denoted as $\left\langle v_{1},v_{2},\ldots,v_{k},v_{1}\right\rangle$ when $k\geq 2$, is a subgraph of $\mathcal{G}$ containing $k$ RDPs $v_{1},\ldots,v_{k}\in\mathcal{R}$ such that there is an arc $(v_{i},v_{i+1})\in\mathcal{A}_{\mathcal{R}}$ for every $i=1,\ldots,k-1$, as well as an arc $(v_{k},v_{1})\in\mathcal{A}_{\mathcal{R}}$. In the presence of compatible RDPs (i.e., when a recipient is compatible with their own paired donor), a cycle can also be of length $k=1$, in which case it consists of a self-loop from an RDP to itself. The length $k$ of a cycle is constrained to be at most $K$. Moreover, a chain of length $\ell$, denoted as $\left\langle v_{0},v_{1},\ldots,v_{\ell-1},\tau\right\rangle$ when $\ell\geq 2$, is a simple path in $\mathcal{G}$, starting with an NDD $v_{0}\in\mathcal{N}$, followed by $\ell-1$ RDPs $v_{1},\ldots,v_{\ell-1}\in\mathcal{R}$, and ending at $\tau$. A chain can also be of length $\ell=1$, in which case it consists of a single arc from an NDD $v_{0}$ to $\tau$. The length $\ell$ of a chain is constrained to be at most $L$. Note that in our definition of $\ell$ and $L$, we include the donation to the terminal node $\tau$. This same convention is used, for example, in Constantino et al., (2013) and Delorme et al., (2024), but in Glorie et al., 2014b , Dickerson et al., (2016) and Mak-Hau, (2017) the final donation is not counted. Whereas our convention is tailored to the case in which $\tau$ represents the DDWL, the other convention is more suited to the case when $\tau$ represents the pool of bridge donors for subsequent matching rounds.
@@
Furthermore, each arc $(u,v)\in\mathcal{A}$ is associated with a weight $w_{uv}$, and the weight of a cycle or chain is defined as the sum of its arcs’ weights. In principle, also the weights $w_{v\tau}$ for $(v,\tau)\in\mathcal{A}_{\mathcal{\tau}}$ can be non-zero, but this might not be desirable when $\tau$ represents the DDWL rather than the pool of bridge donors. It is also possible for a recipient to have multiple donors. One can take such a feature into account by replacing the multiple donors with a single “super-donor” who is considered compatible with a recipient if at least one of the multiple donors is compatible with that recipient. However, this transformation is not applicable in some cases where the weights are donor-specific (see Constantino et al., 2013 for an alternative method).
@@
Given this framework, a feasible solution is defined as a set of exchanges, that is, a collection of vertex-disjoint cycles and chains respecting their respective length limits. The objective of KE-Opt is to find a feasible solution of maximum weight.
@@
We consider two relevant special cases of KE-Opt. First, there is the unweighted version of KE-Opt, where the objective is to maximise the number of transplants. This corresponds to the special case of KE-Opt in which $w_{uv}=1$ for all $(u,v)\in\mathcal{A}$. Second, there is the version of KE-Opt in which there are no NDDs and only cycles are considered. This corresponds to the special case of KE-Opt in which $L=0$. We refer to this as the cycles-only case of KE-Opt, in contrast to the cycles-and-chains case. Unless stated otherwise, we always consider the weighted cycles-and-chains case of KE-Opt.
The following example will serve as a running example throughout this section.
Example 1.
Consider the KE-Opt instance presented in Figure 3. This instance has =$\mathcal{R}=\{1,2,3,4\}$ and $\mathcal{N}=\{5,6\}$, and all arcs have unitary weight (i.e., we consider the unweighted cycles-and-chains case). An optimal solution for $K=2$ and $L=3$ is depicted in bold. This solution consists of cycle $\left\langle 1,4,1\right\rangle$ of length $2$, chain $\left\langle 5,\tau\right\rangle$ of length $1$, and chain $\left\langle 6,2,3,\tau\right\rangle$ of length $3$, resulting in a total of $6$ transplants.
Figure 3: KE-Opt instance with optimal solution for $K=2$ and $L=3$.
[…]

### v2

@@ 2.8 Dataset Generators and Other Software Tools
Dataset generators and software tools play an important role in facilitating simulation studies that help to inform policy decision-making relating to KEPs. Often historical KEP medical data are either not available or not suitable for simulations, making dataset generators highly valuable to researchers and practitioners. Such tools can simulate real-world data under hypothetical changes to the recipient and donor pool, such as a larger number of NDDs and RDPs. These generators fall into two categories: static and dynamic. Static generators produce RDPs and NDDs for a single matching run, whereas dynamic generators generate RDPs and NDDs with arrival and departure times , allowing simulation over a pre-defined time period , and analysis of the long-term evolution of the KEP. Dynamic generators may also generate other temporally relevant information such as periods of absence or scheduling delays. We also review software tools for KEPs. These can allow optimal solutions to be found for a given KEP instance under a range of different constraints and optimality objectives. Moreover, they can allow researchers and practitioners to better understand the long-term effects of different policies informing match runs.
@@
We begin by considering static dataset generators, the first of which was created by Saidman et al., (2006). Their generator takes into account factors such as ABO-compatibility, sensitisation and positive crossmatch probability to create a randomly generated pool of RDPs (note that they did not include NDDs). This generator was used in various subsequent simulations to model KEPs, for instance Roth et al., (2007); Abraham et al., (2007); Dickerson et al., (2013); Constantino et al., (2013); Ashlagi and Roth, (2021). There are 310 instances created by the Saidman generator that are publicly available on PrefLib (Mattei and Walsh, , 2013), initially generated by Dickerson et al., 2012b , some of which also include NDDs.
Delorme et al., (2022) explored differences between datasets produced by the Saidman generator (Saidman et al., , 2006) and historical data from the UK KEP. They found that the Saidman generator typically produces instances that are inconsistent with real UK KEP datasets with respect to a number of measures , such as the density of the compatibility graph. They implemented a new dataset generator to produce instances reflecting key characteristics of UK KEP data; also, their generator can include NDDs.
Two additional static generators were created by Dickerson et al., 2012b and Ashlagi et al., (2013). The former altered the Saidman generator to create sparser instances that reflect the UNOS pool (UNOS being a leading US KEP), whilst the latter created only two types of recipients within their instances: those with low and high PRA [^4] PRA, which stands for Panel Reactive Antibody, measures the sensitisation of a recipient, and indicates the likelihood of the recipient being incompatible with a random blood-group compatible donor from the population.. These generators were used by Arslan et al., (2024) in addition to the aforementioned Saidman instances from PrefLib.
@@
Santos et al., (2017) created a modular and configurable dynamic dataset generator and KEP simulator that allows KEP pools to be created with diverse characteristics, allowing for incompatible RDPs, recipients with multiple donors, compatible RDPs and NDDs. Their tool includes configuration, pool management and optimisation modules and allows the simulation to be controlled by various parameters. The authors also compared their generation and simulation tool with various other dynamic generators in the literature. A more recent dynamic simulator, known as the ENCKEP Simulator, was developed as part of the ENCKEP COST Action (ENCKEP, , 2021). Its optimisation engine allows a range of constraints and optimality objectives to be specified and evaluated in a simulated KEP. It was extended and used in simulations by Matyasi and Biró, (2023) and Druzsin et al., (2024). The former advanced the ENCKEP simulator to test re-optimisation strategies, whilst the latter expanded it to allow for various optimisation criteria and evaluated these criteria in an international setting. Furthermore, Carvalho et al., (2024) used dynamic simulations based on the Canadian Kidney Paired Donation Programme to learn optimal weights to associate with each RDP and NDD used in their selection method. Others have used dynamic dataset generators to model the characteristics of programmes in the USA, such as Sönmez et al., (2020).
@@
Here we mention some additional resources, including web applications and visualisers for kidney exchange, namely (i) a kidney exchange “toolkit” that can solve KE-Opt instances under different hierarchical optimality objectives, written by James Trimble and available from https://www.dcs.gla.ac.uk/~jamest/weighted-toolkit; (ii) a KEP static dataset generator, written by William Pettersson and James Trimble, as described in Delorme et al., (2022) and available from https://wpettersson.github.io/kidney-webapp; and (iii) an interactive kidney exchange game for outreach and public engagement, written by William Pettersson and available from https://www.optimalmatching.com/IP-MATCH/kidney_exchange_game.
@@ 2.9 Emerging Topics
Recent medical advancements have enabled the possibility of transplants between donors and recipients that are ABO- or HLA-incompatible (Andersson and Kratz, , 2020; MacMillan et al., , 2023). However, these transplants can be expensive, resource-intensive and recipients may have to take immunosuppressant medication post-transplant (see, e.g., the overviews from Montgomery et al., 2011 and Colaneri, 2014). Aziz et al., (2021) modelled KE-Opt in the presence of half-compatible arcs, which represent arcs $(v_{1},v_{2})$ in the underlying compatibility graph where the recipient of $v_{2}$ is not immediately compatible with the donor of $v_{1}$, but can become compatible with this donor through a medical procedure or treatment, such as the use of immunosuppressants. They studied the problem of finding an optimal set of exchanges given a budget that limits the number of half-compatible arcs that can be used. The special case of the problem in which all arcs are either compatible or half-compatible was studied by Heo et al., (2021, 2022) from a mechanism design point of view and by Delorme et al., (2025) from a modelling point of view.
@@
As mentioned in Section 1, KEPs may include compatible RDPs, where their participation could give a better match for the recipient compared to their willing donor, and could benefit the remaining KEP pool by giving additional options, especially for hard-to-match recipients (Gentry et al., , 2007). Sönmez and Ünver, (2014) studied Pareto efficient sets of pairwise exchanges in the presence of compatible pairs. Li et al., (2019) presented a new algorithm and conducted simulations for KEPs with compatible RDPs, taking into account the fact that compatible RDPs will typically expect to be matched quickly (or else they would leave the pool and match directly). Sönmez et al., (2020) proposed incentivising compatible RDPs to join a KEP by providing the recipient with a priority in the DDWL, should they require a repeat transplant. Finally, we mention the work of Balbuzanov, (2020), who studied incentivising compatible pairs to participate, in a setting with ordinal recipient preferences, considering criteria such as individual rationality and Pareto optimality.
@@
Another initiative, namely Global Kidney Exchange (GKE), has recently been the focus of some attention in the literature since it was first proposed by Rees et al., (2017). Typically an exchange in GKE will include an RDP with a hard-to-match recipient from a developed country with its own KEP, such as the US, and an RDP from a developing country, where even if this second RDP is a compatible pair, they may be financially incompatible meaning that they cannot afford the cost of a direct donation. This second RDP will travel to the first country, and the surgery and post-transplant care is paid for by the originating KEP in the first country. Although GKE can provide additional transplantation opportunities, the ethics of this form of kidney exchange have been widely debated (see, e.g., Minerva et al., 2019; Ambagtsheer et al., 2020).
Non-Simultaneous Exchanges
We next survey different types of exchanges that relax the requirement for exchanges to be simultaneous. Roth et al., (2006) considered list exchanges, where the donor of an RDP can donate a kidney either to the DDWL, or to another RDP as part of a chain, in exchange for their recipient gaining high priority on the DDWL. They considered incorporating list exchanges and chains triggered by NDDs to KEPs, showing their ability to enhance the success of KEPs. Akbarpour et al., (2025) studied a different form of non-simultaneous exchanges, namely unpaired kidney exchanges. Their algorithm allows a paired recipient to receive a compatible kidney even if their paired donor is not selected (or vice versa). If not selected, the remaining, now unpaired, donor (or recipient) will stay in the pool and donate (receive) a kidney whenever a compatible unpaired recipient (paired or deceased donor) becomes available. While such non-simultaneous exchanges naturally improve the outcomes of KEPs (as the problem is less constrained), it also poses a number of ethical and practical challenges, as highlighted by the authors.
@@ 2.10 Other Related Surveys
There are a range of surveys that emphasise different aspects of KEPs. For example, Mak-Hau, (2017) gave a comprehensive survey of the state of the art in terms of ILP models for kidney exchange at the time of writing. Ashlagi and Roth, (2021) also focussed on summarising KEPs from an operational point of view, specifically in terms of what makes a KEP successful, and they also outlined a range of directions for future research in relation to KEPs. Broad historical overviews of the development of KEPs were given by Gentry et al., (2011), Glorie et al., 2014a and Kher and Jha, (2020), and the role that OR plays has been highlighted in relation to specific KEPs, namely the Alliance for Paired Kidney Donation (Anderson et al., 2015a, ) and the UK Living Kidney Sharing Scheme (Manlove, , 2018). Sönmez and Ünver, (2013); Ashlagi, (2023); Sönmez and Ünver, (2025) survey ed KEPs from a market design perspective , whilst the overview of van Basshuysen, (2020) was given from a philosophical and ethical standpoint. Viana et al., (2023) discussed the complexity of KE-Opt, presenting models and algorithms, and also covered stochastic and robust models , whilst Doval, (2025) provided a detailed survey of dynamic matching, making reference to dynamic KEPs. We also mention the short recent survey given by Sharifi, (2025).
Several papers presented descriptions of the national KEPs in various countries, including Australia (Cantwell et al., , 2015; Sypek et al., , 2017), Canada (Malik and Cole, , 2014), France (Combe et al., , 2019), The Netherlands (de Klerk et al., , 2005), South Korea (Kwak et al., , 1999), Spain (Bofill et al., , 2017), Turkey  (Kutlu-Gündoğdu et al., , 2018), the UK (Johnson et al., , 2008) , and the USA (Agarwal et al., , 2018; Flechner et al., , 2018); and of potential national KEPs for Germany (Ashlagi et al., , 2025) and India (Kute et al., , 2021). When focusing on KEPs across Europe, differences in their medical practices were surveyed by Biró et al., 2019a , while differences in their modelling and optimisation were surveyed by Biró et al., (2021). In addition, Bastos et al., (2024) surveyed KEPs across Latin America and the Caribbean.
@@ 3 A Detailed Exposition of ILP Formulations for KE-Opt
Whilst some algorithms for KE-Opt were surveyed in Section 2.1, we deferred our detailed coverage of ILP-based approaches to this section. We focus on the fundamental version of KE-Opt, as introduced in Section 1, excluding elements such as hierarchical objectives, deceased-donor-initiated chains, and other extensions discussed in Section 2. However, we emphasise that ILP models incorporating these advanced extensions are typically built upon the formulations covered here.
This section is structured as follows. In Section 3.1 we lay the foundations for our formal descriptions of a range of ILP models by providing a formal definition of KE-Opt, together with associated notation and terminology. Next, in Section 3.2 we introduce the ILP models for the version of KE-Opt in which only cycles are considered, after which we explain that these models can be adapted to instead model a version of KE-Opt with only chains. This is followed by a discussion on how to model the cycles-and-chains version of KE-Opt. Subsequently, in Sections 3.3-3.7 we present the ILP models in detail , and in Section 3.8 we compare some theoretical properties of the ILP models.
@@ 3.1 Formal Problem Statement
An instance of KE-Opt involves the set $\mathcal{R}$ of RDPs, the set $\mathcal{N}$ of NDDs and the directed compatibility graph $\mathcal{G}=(\mathcal{V},\mathcal{A})$, as well as the limit $K$ on the maximum cycle length and the limit $L$ on the maximum chain length. Vertex set $\mathcal{V}=\mathcal{R}\cup\mathcal{N}\cup\{\tau\}$ contains a node for every RDP in $\mathcal{R}$ and for every NDD in $\mathcal{N}$, and also a terminal node $\tau$ that represents either the DDWL or the pool of bridge donors for subsequent matching runs. We assume that $\mathcal{R}$ and $\mathcal{N}$ are ordered sets, and unless stated otherwise, we let $\mathcal{R}=\{1,\ldots,|\mathcal{R}|\}$ and $\mathcal{N}=\{|\mathcal{R}|+1,\ldots,|\mathcal{R}|+|\mathcal{N}|\}$. Arc set $\mathcal{A}=\mathcal{A}_{\mathcal{R}}\cup\mathcal{A}_{\mathcal{N}}\cup\mathcal{A}_{\tau}$ contains three types of arcs. First, there is an arc $(u,v)\in\mathcal{A}_{\mathcal{R}}$ for each $u,v\in\mathcal{R}$ such that the donor of RDP $u$ is compatible with the recipient of RDP $v$. Second, there is an arc $(u,v)\in\mathcal{A}_{\mathcal{N}}$ for each $u\in\mathcal{N}$ and $v\in\mathcal{R}$ such that NDD $u$ is compatible with the recipient of RDP $v$. Third, there is an arc $(v,\tau)\in\mathcal{A}_{\tau}$ from every RDP or NDD $v\in\mathcal{R}\cup\mathcal{N}$ to the terminal node $\tau$.
@@
This definition of $\mathcal{G}$ allows for modelling any sequence of exchanges by either a cycle or a chain. A cycle of length $k$, denoted as $\left\langle v_{1},v_{2},\ldots,v_{k},v_{1}\right\rangle$ when $k\geq 2$, is a subgraph of $\mathcal{G}$ containing $k$ distinct RDPs $v_{1},\ldots,v_{k}\in\mathcal{R}$ such that there is an arc $(v_{i},v_{i+1})\in\mathcal{A}_{\mathcal{R}}$ for every $i=1,\ldots,k-1$, as well as an arc $(v_{k},v_{1})\in\mathcal{A}_{\mathcal{R}}$. In the presence of compatible RDPs (i.e., when a recipient is compatible with their own paired donor), a cycle can also be of length $k=1$, in which case it consists of a self-loop from an RDP to itself. The length $k$ of a cycle is constrained to be at most $K$. Moreover, a chain of length $\ell$, denoted as $\left\langle v_{0},v_{1},\ldots,v_{\ell-1},\tau\right\rangle$ when $\ell\geq 2$, is a simple path555 A simple path is a path that contains no repeated vertices. in $\mathcal{G}$, starting with an NDD $v_{0}\in\mathcal{N}$, followed by $\ell-1$ RDPs $v_{1},\ldots,v_{\ell-1}\in\mathcal{R}$, and ending at $\tau$. A chain can also be of length $\ell=1$, in which case it consists of a single arc from an NDD $v_{0}$ to $\tau$. The length $\ell$ of a chain is constrained to be at most $L$. Note that in our definition of $\ell$ and $L$, we include the donation to the terminal node $\tau$. This same convention is used, for example, in Constantino et al., (2013) and Delorme et al., (2024), but in Glorie et al., 2014b , Dickerson et al., (2016) and Mak-Hau, (2017) the final donation is not counted. Whereas our convention is tailored to the case in which $\tau$ represents the DDWL, the other convention is more suited to the case when $\tau$ represents the pool of bridge donors for subsequent matching runs.
@@
Furthermore, each arc $(u,v)\in\mathcal{A}$ is associated with a weight $w_{uv}$, and the weight of a cycle or chain is defined as the sum of its arcs’ weights. In principle, there are situations in which the weights $w_{v\tau}$ for $(v,\tau)\in\mathcal{A}_{\mathcal{\tau}}$ should be set to zero, for example for bridge donors, as the donation from such a donor would potentially be counted in the current matching run as well as in a future matching run. It is also possible for a recipient to have multiple donors. One can take such a feature into account by replacing the multiple donors with a single “super-donor” who is considered compatible with a recipient if at least one of the multiple donors is compatible with that recipient. However, this transformation is not applicable when the weight of an arc $(u,v)$ depends on which of recipient $v$’s donors is used, for example when considering donor-to-donor age difference (see Constantino et al., 2013 for an alternative method).
@@
Given this framework, a feasible solution is defined as a set of exchanges, that is, a collection of vertex-disjoint666 Vertex-disjointness does not apply to $\tau$, as $\tau$ is included in every chain. cycles and chains respecting their respective length limits. The objective of KE-Opt is to find a feasible solution of maximum weight.
@@
We consider two relevant special cases of KE-Opt. First, there is the unweighted version of KE-Opt, where the objective is to maximise the number of transplants. This corresponds to the special case of KE-Opt in which $w_{uv}=1$ for all $(u,v)\in\mathcal{A}$777 Alternatively, one could also set $w_{v\tau}=0$ for all $(v,\tau)\in\mathcal{A}_{\mathcal{\tau}}$ and add the constant $|\mathcal{N}|$ to the objective function.. Second, there is the version of KE-Opt in which there are no NDDs and only cycles are considered. This corresponds to the special case of KE-Opt in which $L=0$. We refer to this as the cycles-only case of KE-Opt, in contrast to the cycles-and-chains case. Unless stated otherwise, we always consider the weighted cycles-and-chains case of KE-Opt.
In Appendix D, we provide an example instance of KE-Opt. We also use that instance as a running example to illustrate the models introduced in the following sections.